WEYL ALGEBRA MODULES

VIKTOR BEKKERT GEORGIA BENKART VYACHESLAV FUTORNY

Abstract. We investigate weight modules for finite and infinite Weyl algebras, classifying all such simple modules. We also study the representation type of the blocks of locally-finite weight module categories and describe indecomposable modules in tame blocks.